rebeccaxhawaii (rebeccaxhawaii):

Let f(x) = 8x^3 - 22x^2 - 4 and g(x) = 4x - 3. Find f(x)/g(x)

OpenStudy (astrophysics):

Divide the two functions

rebeccaxhawaii (rebeccaxhawaii):

yeah obviously but i need help

OpenStudy (astrophysics):

Have you tried factoring

rebeccaxhawaii (rebeccaxhawaii):

well i got 2x^2-4x=13/4x-3 but thats not one of my choices

OpenStudy (astrophysics):

Let me ask you something, are all the signs correct

OpenStudy (astrophysics):

There's not much to simplify here, check if you posted correctly

rebeccaxhawaii (rebeccaxhawaii):

i mean i got 2x^2-4x+13/4x-3

OpenStudy (astrophysics):

\[\frac{ 2(4x^3-11x^2-2) }{ 4x-3 }\]

OpenStudy (astrophysics):

Not much you can do

rebeccaxhawaii (rebeccaxhawaii):

no\[4x-3\sqrt{8x ^{3}-22x ^{2}-4}\]

OpenStudy (astrophysics):

Whattt you didn't even say that..

rebeccaxhawaii (rebeccaxhawaii):

sorry i inserted it wrong the first time

OpenStudy (astrophysics):

Wait so is \[f(x) = \sqrt{8x^3-22x^2-4}\] and \[g(x) = 4x-3\]

rebeccaxhawaii (rebeccaxhawaii):

no that is a division sign

OpenStudy (astrophysics):

\[\frac{ 4x-3 }{ 8x^3-22x^2-4 }\]

rebeccaxhawaii (rebeccaxhawaii):

yes

OpenStudy (astrophysics):

Still not much you can do

rebeccaxhawaii (rebeccaxhawaii):

yes but there is a solution

rebeccaxhawaii (rebeccaxhawaii):

no there is not

OpenStudy (michele_laino):

I think that we can apply the procedure of long division

rebeccaxhawaii (rebeccaxhawaii):

well there is an answer i just messed up in some part and thats why im here

rebeccaxhawaii (rebeccaxhawaii):

yes this is what i need help with the lesson is titled polynomial long division

OpenStudy (michele_laino):

using the procedure of long division I got a different result

OpenStudy (michele_laino):

here is what I got: \[\huge 2{x^2} - 4x - 3 - \frac{{13}}{{4x - 3}}\] is it the right result?

rebeccaxhawaii (rebeccaxhawaii):

YES THAT IS CORRECT thanks so much

rebeccaxhawaii (rebeccaxhawaii):

see guys there is an answer theres no "well theres not much you can do"

OpenStudy (michele_laino):

please try to do such division on your own, since the direct answer is against the Code of Conduct

rebeccaxhawaii (rebeccaxhawaii):

k well thanks anyway

OpenStudy (michele_laino):

:)